Here are simplified financial statements for Watervan Corporation:

INCOME STATEMENT
(Figures in $ millions)
Net sales $

892.00

Cost of goods sold

752.00

Depreciation

42.00

Earnings before interest and taxes (EBIT) $

98.00

Interest expense

23.00

Income before tax $

75.00

Taxes

15.75

Net income $

59.25

BALANCE SHEET
(Figures in $ millions)
End of Year Start of Year
Assets
Current assets $

380

$

334

Long-term assets

280

233

Total assets $

660

$

567

Liabilities and shareholders’ equity
Current liabilities $

205

$

168

Long-term debt

119

132

Shareholders’ equity

336

267

Total liabilities and shareholders’ equity $

660

$

567

The company’s cost of capital is 8.5%.

a. Calculate Watervan’s economic value added (EVA). (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.)

b. What is the company’s return on capital? (Use start-of-year rather than average capital.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

c. What is its return on equity? (Use start-of-year rather than average equity.) (Enter your answer as a percent rounded to 2 decimal places.)

d. Is the company creating value for its shareholders?
Answer #1

(a) EVA = [EbIT x (1 - tax rate)] - (Cost of capital x invested capital)

= 98 millions x (1 - 0.21) - (8.5% x 660 million)

= 77.42 millions - 56.1 millions

= 21.32 millions

working :- Tax rate = 15.75/75 x 100 = 21%

Invested capital is total assets.

(b) Return on capital = Net income/invested capital

= 59.25/567 x 100 = 10.45%

(c) Retuen on equity = net Income/shareholders equity

= 59.25/267 x 100 = 22.19%

(d) Yes, company is creating value for its shareholders because their return on equity is very high.